A single card is drawn from a standard deck of cards. Find the following probabilities. A face card is a jack, queen, or king.
P(jack/red)=

Answer by stanbon(75887)

A single card is drawn from a standard deck of cards. Find the following probabilities. A face card is a jack, queen, or king.
P(jack/red)= 2/52 = 1/26
=========================
Cheers,
Stan H.
=========================